Suppose a meter stick made of steel and one made of invar are the same length at 0°C. What is their difference in length at 30.0 C? Repeat the calculation for two 33.0 m long surveyor's tapes.
Answer #1

Ans:-

Invar 1.2 10^-6 /°C
Steel 13 10^-6 /°C

a) Δ L = α L° Δt
invar 1.2 10^-6 * 1000 * 30 = 0.036 mm
steel 13 10^-6 * 1000 * 30 = 0.39 mm
difference in length at 30.0°C = 0.39 - 0.036 = 0.354 mm

b)
invar 0.036 *33 = 1.188 mm
steel 0.39 * 33 = 12.87 mm
difference in length at 39.0°C = 12.87 -1.188 = 11.682 mm